OpenStudy (anonymous):

Please explain how to simplify this: (x^3+x^2-x-1)/(x^3-x^2-x+1)

OpenStudy (accessdenied):

Do you know how to factor the numerator / denominator? It looks like, if you group the first two terms and the last two terms in each expression, you can factor both by grouping. (x^3 + x^2) + (-x - 1), (x^3 - x^2) + (-x + 1)

OpenStudy (anonymous):

let me try

OpenStudy (accessdenied):

Hmm... were you able to get it?

OpenStudy (accessdenied):

(x^2)(x + 1) + (-1)(x + 1) = (x^2 - 1)(x + 1) (x^2)(x - 1) + (-1)(x - 1) = (x^2 - 1)(x - 1) x^2 - 1 cancels You are left with: (x + 1)/(x - 1)

OpenStudy (anonymous):

I got it thanks